OpenStudy (anonymous):

Ricardo Gutierrez has a box of coins that he uses when playing poker with his friends the box currently contains 49 coins, consisting of pennies, dimes and quarters The number of pennies is equal to the number of dimes and the total value is$5.62 How many of each denomination of coin does he have?

OpenStudy (dumbcow):

p + d+q = 49 p +10d+25q = 562 p = d by substitution: 2d +q = 49 11d +25q = 562 solve system with substitution --> q = 49 -2d 11d +25(49-2d) = 562 solve for d ...